OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5) (Whats the answer thats all I need to know)

OpenStudy (anonymous):

I'll give a medal

OpenStudy (zale101):

\[slope=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Is it -1?

OpenStudy (zale101):

Can you show me how you got it?

OpenStudy (anonymous):

No, can you help me with a few more?

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 (Again just need the answer, I love you guys thanks!)

OpenStudy (zale101):

What did you get for the first problem? For the second problem, you need to know what's the point slope formula. And again, all you have to do is plug and chug with the given point and slope. What is the point slope formula?

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 A. y - 6 = 1/2 (x + 3) B. y = 1/2x - 6 C. y + 3 = -1/5 (x - 2) D. y - 2 = -1/5 (x + 3)

OpenStudy (zale101):

Point slope formula: \(\Large y-y_1=m(x-x_1)\) Where m is the slope 1/2 x1,y1 is the given point (6, -3)

OpenStudy (zale101):

plug all the given info in

OpenStudy (anonymous):

Please all I need is the answer

OpenStudy (anonymous):

Please